### v2

# 1(2), Leninskie gory, Moscow 119991, Russian Federation

Abstract
A method for calculating the $1/d$ expansion coefficients for solutions of integration by parts relations for Feynman integrals is presented. The idea is to use linear substitutions to transform these relations to an explicitly recursive form. The possibility of such a transformation is demonstrated for several families of massless (with one massive line) vacuum integrals up to the 7-loop level.
Keywords:
Higher-Order Perturbative Calculations, Renormalization and Regularization, Scattering Amplitudes
arxiv: 2407.09112
1 Introduction
Multiloop Feynman integrals evaluation is an important ingridient of higher-order perturbative calculations. A very efficient tool in this sence is integration by parts (IBP) relations Chetyrkin:1981qh ; Tkachov:1981wb which are used in a large number of specialized computer programs Gorishnii:1989gt ; Anastasiou:2004vj ; Smirnov:2019qkx ; Klappert:2020nbg ; Lee:2012cn ; Studerus:2009ye ; vonManteuffel:2012np ; Ruijl:2017cxj .
As a result, many record-breaking calculations have been performed; for a few recent examples, see Georgoudis:2021onj ; Chakraborty:2022yan ; Lee:2023dtc ; Behring:2020cqi ; Bern:2023ccb . However, extension of the IBP method to higher perturbation orders faces significant challenges such as growing computer and hand-working demands, increasing the number of subcases that need to be considered and therefore the probability of errors of various origins. Although some reserve still exists, both purely mathematical and algorithmic improvements are highly desirable for further progress.
The $1/d$ expansion (where $d$ is the dimensional regularization parameter) in the context of applying the IBP relations was proposed in Baikov:2005nv ; Baikov:2007zza and was successfully used in the number of calculations (see, for example Baikov:2008jh ; Baikov:2009bg ; Baikov:2010je ; Baikov:2012er ; Baikov:2012zm ; Baikov:2012zn ; Baikov:2014qja ; Baikov:2016tgj ; Baikov:2017ujl ). From a technical point of view this requires expansion in the limit $1/d\rightarrow 0$ of integrals with integrands depending on $P(x_{i})^{d/2}$ where $P(x_{i})$ is some multivariate polinomial. This expansion causes the size of intermediate expressions to rapidly increase and calculations to a sufficiently high order of $1/d$ become possible only after manual optimization (smart redefinition of variables, etc.), which was carried out individually for different subclasses of integrals. Bearing in mind the significant increase in the subclasses number at the next loop(s) level, in this article we reformulate the problem as constructing solutions to IBP relations in the form of formal series of $1/d$ and propose linear substitutions that transform these relations into an explicitly recursive form.
The plan of the paper is the following. In section 2 we fix the notations and breafly recall the main points of the Baikov:1996rk ; Baikov:1996iu ; Baikov:2005nv ; Baikov:2007zza . Then in section 3 we introduce (for one-parameter case) specific linear substitutions, which we apply in section 4 to the IBP relations. In the next section 5 we consider several examples of massless (with one massive line) vacuum integral families up to the 7-loop level. Finally, in section 6 we discuss a possible extension to more complex cases.
2 Integration by parts relations for Feynman integrals
Let us consider the $d$-dimensional $L$-loop vacuum Feynman integrals:
$F(n_{1},\ldots,n_{N};d;m^{2})=\int\ldots\int\frac{\mbox{d}^{d}p_{1}\ldots\mbox{d}^{d}p_{L}}{D_{1}^{n_{1}}\ldots D_{N}^{n_{N}}}\equiv(m^{2})^{Ld/2-\Sigma n_{i}}f(n_{1},\ldots,n_{N};d),$ (1)
where $p_{i}$ (${i=1,\ldots,L}$) are loop momenta and
$D_{a}$ $=$ $A^{ij}_{a}p_{i}\cdot p_{j}-\mu_{a}m^{2},\quad a=1,\ldots,N=L(L+1)/2$ (2)
are inverse propagators. The matrix $A^{ij}_{a}$ is supposed to be symmetrical with respect to upper indices. The $1/D_{a}$ can be raised as to positive powers (“denominators” or “lines”) as non-positive (“numerators”). Summation over repeating indices is implied and in many cases (unless it’s confusing) they will be omitted. In what follows we will set $m^{2}=1$; we will also use condensed notation $(n_{1},\ldots,n_{N};d)\equiv(\underline{n})$.
We assume that the set (2) is complete and we can express scalar products $p_{i}\cdot p_{j}$ through $D_{a}$:
$p_{i}\cdot p_{k}$ $=$ $\hat{A}^{ik}_{a}(D_{a}+\mu_{a})\,.$ (3)
The matrix $\hat{A}^{ij}_{a}$ is also symmetrical with respect to $(ij)$ and is inverse to $A^{kl}_{a}$ in the sence that
$\hat{A}^{ij}_{a}A^{kl}_{a}$ $=$ $(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})/2,\quad A^{ik}_{a}\hat{A}^{ik}_{b}=\delta_{ab}.$ (4)
Please note that in Baikov:2005nv for the coefficients in (3) we used the notation $\tilde{A}$; we change it to $\hat{A}$ because we want to reserve the $\sim$” sign for “soft” objects (see section 4).
Lets define
${\bf I}^{-}_{a}f(\ldots,n_{a},\ldots)$$\equiv$ $f(\ldots,n_{a}-1,\ldots),$
${\bf I}^{+}_{a}f(\ldots,n_{a},\ldots)$$\equiv$ $n_{a}f(\ldots,n_{a}+1,\ldots)$
(no summation in $n_{a}$ in the last relation).
IBP relations can be obtained Chetyrkin:1981qh ; Tkachov:1981wb by inserting the operator $(\partial/\partial p_{i})\cdot p_{k}$ in the integrand, equating the integral of the derivative to zero and expressing all the terms resulting from the differentiation through the integrals with shifted $(\underline{n})$. Starting from eq. (1) we obtain (upto $1/2$ factor)
$0=R_{ik}(\underline{n})$ $=$ $\big{(}(d/2)\,\delta_{ik}-\hat{A}^{kl}_{a}({\bf I}^{-}_{a}+\mu_{a})A_{b}^{il}{\bf I}_{b}^{+}\big{)}\,f(\underline{n})\,,$ (5)
$=$ $\big{(}(d-L-1)/2\,\delta_{ik}-A_{b}^{il}{\bf I}_{b}^{+}\,\hat{A}^{kl}_{a}({\bf I}^{-}_{a}+\mu_{a})\big{)}f(\underline{n})\,,$ (6)
where we use $[{\bf I}^{-}_{a},{\bf I}^{+}_{b}]=\delta_{ab}$ to derive eq. (6).
A linear combination of $R_{ik}(\underline{n})$ with shifted $(\underline{n})$ corresponds to shifts of $(\underline{n})$ in $f(\underline{n})$ and hence to the insertion of some combination of $({\bf I}^{-}_{a},{\bf I}^{+}_{a})$ just before $f(\underline{n})$.
In many practically important cases eq. (5) can be used to relate integrals with arbitrary values of $(\underline{n})$ to a linear combination of a finite number of irreducible (master) integrals:
$f(\underline{n})$ $=$ $\sum_{k}c_{k}(\underline{n})\,M_{k},\quad M_{k}=f(\underline{n}_{k}),$ (7)
where $f(\underline{n}_{k})$ are chosen as the most convinient for direct calculation, usually with values $(\underline{n}_{k})$ from the set $(-1,0,+1)$.
Such a reduction can be achieved as by deriving relations with an explicite recursive structure (MINCER Gorishnii:1989gt for 3-loop propagator integrals, Forcer Ruijl:2017cxj for the next loop level) as by Laporta approach Laporta:2000dsw with a number of implementations Anastasiou:2004vj ; Smirnov:2019qkx ; Klappert:2020nbg ; Lee:2012cn ; Studerus:2009ye ; vonManteuffel:2012np . In any case, reduction to irreducible integrals during actual calculations requires huge computer resources, so any progress in this field is very desirable. In this context, a direct calculation of c(n) was proposed Baikov:1996rk ; Baikov:1996iu . Indeed, if $M_{k}$ are linear independent, then $c_{k}(\underline{n})$ are rational in $d$ solutions of IBP relations and can be obtained from any other solutions set as linear combinations satisfying the condition $c_{i}(\underline{n}_{k})=\delta_{ik}$.
There are natural boundary conditions on $c_{k}(\underline{n})$, which follow from the assumption that no additional denominators (lines) appear on the right side of eq. (7) (since the purpose of the reduction procedure is to relate the original integral with simpler ones):
$c_{k}(\underline{n})$$\neq$ 0only $0\quad\mbox{only if}\,\,n_{a}>0\,\,\mbox{for all}\,\,a\in H_{k},$
where $H_{k}=\{a_{1},\dots,a_{k}\}$ is the set of line numbers of the irreducible integral $M_{k}$. We will use also (for auxiliary constructions) an alternative boundary condition for $a\notin H_{k}$:
$f_{k}(\underline{n})$$\neq$ 0only $0\quad\mbox{only if}\,\,n_{a}<1\,\,\mbox{for all}\,\,a\notin H_{k}.$ (8)
Here and below we consider $f(\underline{n})$ as some solution of recurrence relations (and integrals (1) as a particular example).
In Baikov:1996rk ; Baikov:1996iu $c_{k}(\underline{n})$ were constructed using the auxiliary integral representation
$f(\underline{n})$ $=$ $\int dx_{1}\dots dx_{N}\,f(\underline{x})/(x_{1}^{n_{1}}..\ x_{N}^{n_{N}}),$ (9)
where $f(\underline{x})$ should be chosen according to recurrence relations; for example, relations for integrals (1) require
$f(\underline{x})=\det(\hat{A}^{ik}_{a}(x_{a}+\mu_{a}))^{(d-L-1)/2}.$ (10)
Integrals (9) can be expanded in the $1/d\rightarrow 0$ limit Baikov:2005nv ; Baikov:2007zza but as mentioned in the introduction this is difficult to extend to the next loop(s) level. Therefore, as a more promising alternative, we are going to explore ways to find the solutions to IBP relations as formal series in $1/d$.
3 Linear substitutions in recurrence relations: one-parameter case
Let’s consider the one-parameter recurrence relation:
00 0 $=$ $R(n)=R({\bf I}_{f}^{-},{\bf I}_{f}^{+})f(n),$ (11)
where by definition
${\bf I}_{f}^{+}f(n)$$\equiv$ $nf(n+1),\quad{\bf I}_{f}^{-}f(n)\equiv f(n-1),$ (12)
and hence
$\big{[}{\bf I}_{f}^{-},{\bf I}_{f}^{+}\big{]}$$\equiv$ ${\bf I}_{f}^{-}{\bf I}_{f}^{+}-{\bf I}_{f}^{+}{\bf I}_{f}^{-}=1,$
$\big{[}{\bf I}_{f}^{-},({\bf I}_{f}^{+})^{k}\big{]}$ $=$ $k\,({\bf I}_{f}^{+})^{k-1},\quad\big{[}{\bf I}_{f}^{+},({\bf I}_{f}^{-})^{k}\big{]}=-k\,({\bf I}_{f}^{-})^{k-1},$
$\big{[}{\bf I}_{f}^{-},S({\bf I}_{f}^{+})\big{]}$ $=$ $\partial_{{\bf I}_{f}^{+}}S({\bf I}_{f}^{+}),\quad\big{[}{\bf I}_{f}^{+},S({\bf I}_{f}^{-})\big{]}=-\partial_{{\bf I}_{f}^{-}}S({\bf I}_{f}^{-}),$
where $S({\bf I})$ is any formal seria in ${\bf I}$.
We are going to study the transformation of eq. (11) after linear redefinition of $f(n)$. The idea is to replace $f(n)$ with some combination of $g(n)$ to obtain a simpler recurrence relation for $g(n)$, solve them and re-express the results in terms of $f(n)$.
Let’s start with eq. (11) with the boundary condition
$f(n<1)=0.$ (13)
Due to eq. (13) if $k>n-1$ then $({\bf I}_{f}^{-})^{k}f(n)=f(n-k)=0$, so
$g(n)$ $=$ $e^{\lambda{\bf I}_{f}^{-}}f(n)\equiv\sum_{k=0}^{\infty}\frac{(\lambda{\bf I}_{f}^{-})^{k}}{k!}f(n)$ (14)
will be equal to the finite linear combination of $f(k)$ with $k=1,\dots,n$. We are going to derive the recurrence relation for $g(n)$, which follows from eq. (11). To do this we need to find the relations between operators ${\bf I}_{f}^{\pm}$ defined by eq. (12) and ${\bf I}_{g}^{\pm}$, acting in the same way on $g(n)$. Let’s start with ${\bf I}_{g}^{+}$:
${\bf I}_{g}^{+}g(n)$$\equiv$ $ng(n+1)=ne^{\lambda{\bf I}_{f}^{-}}f(n+1)=e^{\lambda{\bf I}_{f}^{-}}nf(n+1)=e^{\lambda{\bf I}_{f}^{-}}{\bf I}_{f}^{+}f(n)$
$=$ $\big{(}\big{[}e^{\lambda{\bf I}_{f}^{-}},{\bf I}_{f}^{+}\big{]}+{\bf I}_{f}^{+}e^{\lambda{\bf I}_{f}^{-}}\big{)}f(n)=\big{(}\partial_{{\bf I}_{f}^{-}}(e^{\lambda{\bf I}_{f}^{-}})+{\bf I}_{f}^{+}e^{\lambda{\bf I}_{f}^{-}}\big{)}f(n)=(\lambda+{\bf I}_{f}^{+})e^{\lambda{\bf I}_{f}^{-}}f(n)$
$=$ $({\bf I}_{f}^{+}+\lambda)g(n).$
The relation for ${\bf I}_{g}^{-}$ is more simple:
${\bf I}_{g}^{-}g(n)$$\equiv$ $g(n-1)=e^{\lambda{\bf I}_{f}^{-}}f(n-1)=e^{\lambda{\bf I}_{f}^{-}}{\bf I}_{f}^{-}f(n)={\bf I}_{f}^{-}e^{\lambda{\bf I}_{f}^{-}}f(n)$
$=$ ${\bf I}_{f}^{-}g(n).$
Now we can invert eq. (14):
$f(n)$ $=$ $e^{-\lambda{\bf I}_{f}^{-}}g(n)$ (15)
$=$ $e^{-\lambda{\bf I}_{g}^{-}}g(n).$
Again, due to eqs. (1314) it turns out that $g(n<1)=0$ and eq. (15) consists of a finite number of terms.
The recurrence relation for $g(n)$ follows from the linear combination of eqs. (11) with shifted $n$:
00 0 $=$ $e^{\lambda{\bf I}_{R}^{-}}R(n)=R({\bf I}_{f}^{-},{\bf I}_{f}^{+})e^{\lambda{\bf I}_{f}^{-}}f(n)=R({\bf I}_{f}^{-},{\bf I}_{f}^{+})g(n)$
$=$ $R({\bf I}_{g}^{-},{\bf I}_{g}^{+}-\lambda)g(n).$
The case with the boundary condition
$f(n>0)=0$ (16)
is considered similarly, we present the final formulas:
$g(n)$ $=$ $e^{\lambda{\bf I}_{f}^{+}}f(n),\quad g(n>0)=0,$
$f(n)$ $=$ $e^{-\lambda{\bf I}_{g}^{+}}g(n),$ (17)
${\bf I}_{g}^{-}$ $=$ ${\bf I}_{f}^{-}-\lambda,\quad{\bf I}_{g}^{+}={\bf I}_{f}^{+},$
00 0 $=$ $R({\bf I}_{g}^{-}+\lambda,{\bf I}_{g}^{+})g(n).$ (18)
The boundary condition (16) corresponds to the condition (8), that is, the case when the master integral $M_{k}$ does not have the line $D_{a}$. However, there may be such a line on the left side of eq. (7), which means that $n_{a}$ in $c_{k}(\underline{n})$ can be positive. So, let us extend the substitution (17) to positive $n$. For $n<1$ we can rewrite $g(n)=({\bf I}_{g}^{-})^{-n}g(0)$ and then eq. (17) as follows:
$f(n)$ $=$ $e^{-\lambda{\bf I}_{g}^{+}}({\bf I}_{g}^{-})^{-n}g(0)=({\bf I}_{g}^{-}+\lambda)e^{-\lambda{\bf I}_{g}^{+}}({\bf I}_{g}^{-})^{-n-1}g(0)=\dots=({\bf I}_{g}^{-}+\lambda)^{-n}e^{-\lambda{\bf I}_{g}^{+}}g(0)$ (19)
$=$ $({\bf I}_{g}^{-}+\lambda)^{-n}g(0).$
For $f(n<1)$ the binomial expansion reproduces the finite sum (17) as it should. For $f(n>0)$ we expand the right side of eq. (19) into a formal series in ${\bf I}_{g}^{-}$. From the negative binomial series we get
$f(n>0)$ $=$ $({\bf I}_{g}^{-}+\lambda)^{-n}g(0)$ (20)
$=$ $\sum_{k=0}^{\infty}(-1)^{k}\frac{(n+k-1)!}{(n-1)!\,k!}\lambda^{-n-k}({\bf I}_{g}^{-})^{k}g(0)$
$=$ $\sum_{k=0}^{\infty}(-1)^{k}\frac{(n+k-1)!}{(n-1)!\,k!}\lambda^{-n-k}g(-k),$
that is the infinite sum of $g(n<1)$. However, as we will see in section 4 one can choose $\lambda$ so that the $k$-term in the sum (20) has leading asymptotics $d^{-k/2}$ (see eq. (21)), therefore, by keeping enough terms in this sum and assuming that $g(n<1)$ satisfies eq. (18) we can satisfy (11) in any desired order in $1/d$. As a result, $f(n)$ can be parameterized using $g(k<1)$ for any value of $n$. In what follows, by the expression (17) in the case (16) we will mean the extension (20) for $n>0$ (in the $\lambda=0$ case when (20) is not defined we will set $f(n)=g(n)$).
Finally, if we rescale $f(n)=\alpha^{n}g(n)$, then eq. (11) transforms into
00 0 $=$ $R({\alpha^{-1}\bf I}_{g}^{-},\alpha{\bf I}_{g}^{+})g(n).$
4 Linear substitutions in recurrence relations: Feynman integrals
Let us now return to the general case and construct, if possible, a solution to eq. (5) for some set of propagator indices $H_{k}=\{a_{1},\dots,a_{k}\}$. We will call the index $a\in H_{k}$ “hard”. (substitutions (15) will be applied), and $a\notin H_{k}$ “soft’ (substitutions (1720)).
For any indexed object $x_{a}$ we define projections onto the “hard” and “soft” subspaces: ${\overline{x}}_{a}\equiv x_{\overline{a}}=$ (if $a\in H_{k}$ then $x_{a}$ else 00) and ${\tilde{x}}_{a}\equiv x_{\tilde{a}}=x_{a}-{\overline{x}}_{a}$; as the result $x_{a}=\overline{x}_{a}+\tilde{x}_{a}$. For $x_{a}={\bf I}_{a}^{\pm}$ we have ${\bf I}_{a}^{\pm}={\tilde{\bf I}}_{a}^{\pm}+{\overline{\bf I}}_{a}^{\pm},\,\big{[}{\tilde{\bf I}}_{a}^{\pm},{\overline{\bf I}}_{b}^{\pm}\big{]}=0.$
Let us redefine
$d/2$ $=$ $1/\alpha^{2},\quad\alpha\rightarrow 0$
According to section 3 the substitutions
$f(\underline{n})$ $=$ $\exp(\alpha^{-2}\overline{\lambda}\,\overline{{\bf I}}_{g}^{-}-\tilde{\lambda}\tilde{{\bf I}}_{g}^{+})f^{\prime}(\underline{n}),$
$f^{\prime}(\underline{n})$ $=$ $\alpha^{-\sum\underline{n}}g(\underline{n})$ (21)
will change $\overline{{\bf I}}_{f,a}^{+}=\alpha^{-2}(\alpha\overline{{\bf I}}_{g,a}^{+}+\overline{\lambda}_{a}),\,\tilde{{\bf I}}_{f,a}^{-}=\alpha\tilde{{\bf I}}_{g,a}^{-}+\tilde{\lambda}_{a}$ and therefore eq. (5) will be transformed into
00 0 $=$ $\Big{(}\delta_{ik}-\hat{A}^{kl}(\alpha\,{\bf I}^{-}_{g}+\tilde{\lambda}+\mu)A^{li}\big{(}{\alpha\,\bf I}_{g}^{+}+\overline{\lambda}\big{)}\Big{)}\,g(\underline{n})\,$ (22)
$=$ $\Big{(}\delta_{ik}-\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}\,\overline{\lambda}$
$-\alpha\,\big{(}\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}{\bf I}_{g}^{+}+\hat{A}^{kl}{\bf I}^{-}_{g}A^{li}\overline{\lambda}\big{)}-\alpha^{2}\hat{A}^{kl}{\bf I}^{-}_{g}A^{li}{\bf I}_{g}^{+}\Big{)}g(\underline{n}).$
Compared to eq. (5), we have changed the order of the indices of the matrix $A^{kl}_{a}$ (since it is symmetric). Eqs. (22) generate relations for the expansion coefficients $g_{j}(\underline{n})$ in $g(\underline{n})=\sum_{j=0}^{\infty}\alpha^{j}\ g_{j}(\underline{n})$ (in what follows we omit the subscript $g$ in ${\bf I}^{\pm}_{g}$):
00 0 $=$ $\big{(}\delta_{ik}-\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}\overline{\lambda}\big{)}g_{j}(\underline{n})$ (23)
$-\big{(}\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}{\bf I}^{+}+\hat{A}^{kl}{\bf I}^{-}A^{li}\overline{\lambda}\big{)}g_{j-1}(\underline{n})$
$-\hat{A}^{kl}{\bf I}^{-}A^{li}{\bf I}^{+}g_{j-2}(\underline{n}).$
Equations for $g_{0}(\underline{n})$ do not contain ${\bf I}_{a}^{\pm}$ operators and in general leads to $g_{0}(\underline{n})=0$ (and as a result $g_{j}(\underline{n})=0$), therefore, to obtain non-trivial solutions it is necessary to choose $\lambda_{a}$ satisfies the equation
$\delta_{ik}$ $=$ $\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}\,\overline{\lambda}.$ (24)
Eq. (23) with the condition (24) becomes simpler:
00 0 $=$ $\big{(}\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}{\bf I}^{+}+\hat{A}^{kl}{\bf I}^{-}A^{li}\overline{\lambda}\big{)}g_{j}(\underline{n})$ (25)
$+\hat{A}^{kl}{\bf I}^{-}A^{li}{\bf I}^{+}g_{j-1}(\underline{n}).$
We are going to use eq. (25) to reduce $g_{j}(\underline{n})$ (up to $g_{j-1}(\underline{n})$ terms) to “corner” values: “hard” $\overline{n}_{a}\to 1$ and “soft” $\tilde{n}_{a}\to 0$ (assuming “soft” $\tilde{\lambda}_{a}\neq 0$, see the end of section 3). This would be possible if the $g_{j}(\underline{n})$ part of eq. (25) could be solved explicitly with respect to the operators ($\overline{\bf I}_{a}^{+},\tilde{\bf I}_{a}^{-}$), increasing the distance from the corner. Let us multiply eqs. (25) by $(\hat{A}_{a}A\overline{\lambda})^{ik}$:
00 0 $=$ $\Big{(}\mbox{Tr}\big{(}\hat{A}_{a}A\overline{\lambda}\,\hat{A}(\tilde{\lambda}+\mu)A{\bf I}^{+}\big{)}+\mbox{Tr}\big{(}\hat{A}_{a}A\overline{\lambda}\,\hat{A}{\bf I}^{-}A\overline{\lambda}\big{)}\Big{)}g_{j}(\underline{n})+\mbox{Tr}(\dots)g_{j-1}(\underline{n})$ (26)
$=$ $\big{(}{\bf I}_{a}^{+}+\mbox{Tr}(\hat{A}_{a}A\overline{\lambda}\,\hat{A}_{b}A\overline{\lambda}){\bf I}_{b}^{-}\big{)}g_{j}(\underline{n})+\mbox{Tr}(\hat{A}_{a}A\overline{\lambda}\,\hat{A}{\bf I}^{-}A{\bf I}^{+})g_{j-1}(\underline{n}).$
Here $\mbox{Tr}(\cdot)$ means a contraction of outer indices $(ik)$; we also use the condition (24) and $\mbox{Tr}(\hat{A}_{a}A_{b})=\delta_{ab}$. In the case of “soft” $\tilde{a}$, the $g_{j}(\underline{n})$ part of eq. (26) depends on increasing $\tilde{\bf I}_{a}^{-}$ and decreasing $(\tilde{\bf I}_{b}^{+},\,\overline{\bf I}_{b}^{-})$ operators and if
$\det_{\tilde{a}\tilde{b}}\mbox{Tr}(\hat{A}_{\tilde{a}}A\overline{\lambda}\,\hat{A}_{\tilde{b}}A\overline{\lambda})$$\neq$ $0,$ (27)
where both indices $a,b$ are “soft”, we can express increasing “soft” operators $\tilde{\bf I}_{a}^{-}$ in terms of decreasing ones:
$\tilde{\bf I}_{a}^{-}$$\rightarrow$ $(\tilde{\bf I}_{b}^{+},\,\overline{\bf I}_{b}^{-}).$ (28)
For “hard” $\overline{a}$ eq. (26) explicitly expresses the remaining increasing operators $\overline{\bf I}_{a}^{+}$ in terms of others $(\tilde{\bf I}_{b}^{-},\,\tilde{\bf I}_{b}^{+},\,\overline{\bf I}_{b}^{-})$ and due to eq. (28) in terms of decreasing $(\tilde{\bf I}_{b}^{+},\,\overline{\bf I}_{b}^{-})$ only, so as the result we have an explicite reduction to “corner” values
$(\tilde{\bf I}_{a}^{-},\,\overline{\bf I}_{a}^{+})$$\rightarrow$ $(\tilde{\bf I}_{b}^{+},\,\overline{\bf I}_{b}^{-}).$ (29)
The $L(L+1)/2$ equations (29) are linear combinations of the $L^{2}$ eqs. (26) (and hence of the $L^{2}$ eqs. (25)), so we can expect additional linear relations on ${\bf I}_{a}^{\pm}$, which can lead to a contradiction. Fortunately if (27) and hence (29) are valid this is not the case. Indeed, eqs. (26) are the combinations of eqs. (29). Then the $g_{j}(\underline{n})$ parts of eqs. (25) are the linear combinations of the $g_{j}(\underline{n})$ parts of eqs. (26) with the coefficients $(\hat{A}(\tilde{\lambda}+\mu)A_{a})^{ki}$, so substituting (29) into eq. (25) nullifies this part and no additional relations appear.
The case when some of the “soft” $\tilde{\lambda}_{a}=0$ is more complicated, since the corresponding $\tilde{n}_{a}$ can be both negative and positive and finding out the possibility of reducing them to a minimum value requires a more detailed study of the $g_{j}(\underline{n})$ parts of eqs. (25).
Eqs. (24) leads to several useful consequences. First, $\mbox{Tr}(\cdot)$ of eqs. (24) due to eq. (4) leads to
$L$ $=$ $\overline{\mu}_{a}\,\overline{\lambda}_{a}.$ (30)
Then, $\det(\cdot)$ of eqs. (24) leads to
11 11 $=$ $\det\big{(}\hat{A}(\tilde{\lambda}+\mu)\big{)}\det(A\overline{\lambda})$ (31)
$\Rightarrow$ $\det\big{(}\hat{A}(\tilde{\lambda}+\mu)\big{)}\neq 0.$
Next, $\hat{A}(\tilde{\lambda}+\mu)$ and $A\overline{\lambda}$ considered as matrices with indices $(ik)$ are reciprocal
$A\overline{\lambda}$ $=$ $\hat{A}(\tilde{\lambda}+\mu)^{-1}.$ (32)
Finally, let us multiply eqs. (24) by $\big{(}\hat{A}_{a}\hat{A}(\tilde{\lambda}+\mu)^{\mbox{\small Adj}}\big{)}^{ik}$:
$\partial_{a}\det\big{(}\hat{A}(\tilde{\lambda}+\mu)\big{)}$ $=$ $\overline{\lambda}_{a}\,\det\big{(}\hat{A}(\tilde{\lambda}+\mu)\big{)},$ (33)
where $A^{\mbox{\small Adj}}\equiv\det(A)\,A^{-1}$, $\partial_{a}\det(A\lambda)\equiv\partial_{x_{a}}\det(A^{ik}_{a}x_{a})|_{x_{a}=\lambda_{a}}$. For “soft” $\tilde{a}$ eq. (33) gives
$\partial_{\tilde{a}}\det(\hat{A}(\tilde{\lambda}+\mu))$ $=$ $0$ (34)
Using eq. (32) we can reexpress eq. (27) via $\tilde{\lambda}_{a}$:
00 0 $\neq$ $\det_{\tilde{a}\tilde{b}}\mbox{Tr}\big{(}\hat{A}_{\tilde{a}}A\overline{\lambda}\,\hat{A}_{\tilde{b}}A\overline{\lambda}\big{)}$ (35)
$=$ $\det_{\tilde{a}\tilde{b}}\Big{(}\mbox{Tr}\big{(}\hat{A}_{\tilde{a}}\hat{A}(\tilde{\lambda}+\mu)^{-1}\hat{A}_{\tilde{b}}\hat{A}(\tilde{\lambda}+\mu)^{-1}\big{)}\Big{)}$
$\propto$ $\det_{\tilde{a}\tilde{b}}\Big{(}\partial_{\tilde{a}}\partial_{\tilde{b}}\det\big{(}\hat{A}(\tilde{\lambda}+\mu)\big{)}\Big{)}/\det\big{(}\hat{A}(\tilde{\lambda}+\mu)\big{)},$
where in the third line we use eq. (34) and standard matrix formulas
$\partial_{x}\det(A)$ $=$ $\det(A)\,\mbox{Tr}\big{(}A^{-1}\partial_{x}A\big{)},\quad\partial_{x}(A^{-1})=-A^{-1}\,\partial_{x}(A)\,A^{-1}.$
Please note that eqs. (34, 31, 35) correspond to eqs. (16) from Baikov:2005nv , that is to the conditions for the existence of the $1/d$ expansion of the integral representation (9, 10).
5 Examples
Calculating the parameters $\lambda_{a}$ is the key point for obtaining recursive relations for coefficients of $1/d$ series. To find $\lambda_{a}$, it is necessary to solve eqs. (24), that is, a system of quadratic equations for $L(L+1)/2$ variables (for example, 28 variables for the 7-loop case). However, the bilinear structure of eqs. (24) allows us to significantly reduce this number. First, using the linear equation (30) we can exclude one of $\lambda_{a}$. Moreover, we can obtain additional linear relations on $\lambda_{a}$ checking general combinations of eqs. (24) with polynomials in $\lambda_{a}$ as coefficients. If such relations are found, with their help we can eliminate some $\lambda_{a}$ from eqs. (24) and repeat the same steps for the resulting equations.
Figure 1 presents “hard” $\overline{\lambda}_{a}$ (“soft” $\tilde{\lambda}_{a}$ can be recovered using eq. (32)) for leading sectors of some integral families upto the 7-loop order. Polynomials of the first degree as coefficients were sufficient to reduce the original system of equations to an equation for one variable. [^1] Symbolic manipulations were performed using REDUCE (https://reduce-algebra.sourceforge.io/).
$L=5$$c$$e$$e$$c$$e$$e$$b$$b$$a$$b$$b$$c$
$L=6$$c$$e$$e$$e$$c$$e$$e$$e$$b$$b$$a$$b$$b$$c$$c$
$L=7$$c$$e$$e$$e$$e$$c$$e$$e$$e$$e$$b$$b$$a$$b$$b$$c$$c$$c$
$L=6$$e$$e$$e$$e$$e$$e$$c$$c$$c$$c$$b$$b$$a$$b$$b$
$L=7$$e$$e$$e$$e$$e$$e$$e$$e$$c$$c$$c$$c$$c$$b$$b$$a$$b$$b$
Figure 1: $\overline{\lambda}_{a}$ (roots of eqs. (24)) for some vacuum diagrams. Ordinary lines are massless, double line is massive; $a_{L}=L$, $b_{L}=-\frac{L-2}{2}e_{L}$, $c_{L}=\frac{L(L-2)}{2}e_{L}-L(L-1)$;
for planar diagrams (first row) $e_{5}=\frac{20}{9}$, $e_{6}=\frac{15}{8}$, $e_{7}=(\pm 28\,(61)^{1/2}+1232)/625$;
for nonplanar (second row) $e_{6}=(\pm 3\,(41)^{1/2}+57)/16$, $e_{7}=(\pm 168\,(26)^{1/2}+2268)/625$.
The results show that in 5-, 6-loop planar examples there is at least one $1/d$ solution, in 7-loop planar and 6-, 7-loop nonplanar cases there are at least two. The $L$-dependence of the results was obtained empirically, but there are some indications of its validity for an arbitrary order, we hope to clarify this point in future publications.
Please note that eq. (16) from Baikov:2005nv (eq. (34) in the notation of this article) at the 7-loop level leads to polynomial equations of the 6th degree for 10 variables, which is a much more difficult task.
6 Conclusion
In this paper we present a method to construct solutions of IBP relations for Feynman integrals as formal series in the limit $1/d\rightarrow 0$. The idea is to use substitutions of the form (1517) which transform IBP relations into an explicit recursive form. From a technical point of view, it is necessary to calculate the constants $\lambda$ as solutions to a system of bilinear quadratic equations (24). Although this is relatively simple (and therefore can be automated) in the case where one line is massive and the others are massless, more general kinematic configurations may be more difficult to study. In particular, it may happen that the number of $1/d$ solutions constructed will be less than the number of master integrals and, therefore, not enough to calculate all $c_{i}(\underline{n})$ coefficients in eq. (7). However, even in this case they provide some linear relations between $c_{i}(\underline{n})$ that can be used (in combination with other data) for their reconstruction or at least for verification.
Please note that although in this article we are dealing with recurrence relations for vacuum integrals, the substitutions (15, 17) can also be applied to the case with external momenta; we leave this for future research.