### v3

# 1(2), Leninskie gory, Moscow 119991, Russian Federation

Abstract
A method for calculating the $1/d$ expansion coefficients for solutions of integration by parts relations for Feynman integrals is presented. The idea is to use linear substitutions to transform these relations to an explicitly recursive form. A possible type of such substitutions is proposed for the case of vacuum integrals. Its applicability is shown for several families of massless (with one massive line) vacuum integrals up to the 7-loop level.
Keywords:
Higher-Order Perturbative Calculations, Renormalization and Regularization, Scattering Amplitudes
arxiv: 2407.09112
1 Introduction
Multiloop Feynman integrals evaluation is an important ingredient of higher-order perturbative calculations. A very efficient tool in this sense is integration by parts (IBP) relations Chetyrkin:1981qh ; Tkachov:1981wb which are used in a large number of specialized computer programs Gorishnii:1989gt ; Anastasiou:2004vj ; Smirnov:2019qkx ; Klappert:2020nbg ; Lee:2012cn ; Studerus:2009ye ; vonManteuffel:2012np ; Ruijl:2017cxj .
As a result, many record-breaking calculations have been performed; for a few recent examples, see Georgoudis:2021onj ; Chakraborty:2022yan ; Lee:2023dtc ; Behring:2020cqi ; Bern:2023ccb . However, extension of the IBP method to higher perturbation orders faces significant challenges such as growing computer and hand-working demands, increasing the number of subcases that need to be considered and therefore the probability of errors of various origins. Although some reserve still exists, both purely mathematical and algorithmic improvements are highly desirable for further progress.
The $1/d$ expansion (where $d$ is the dimensional regularization parameter) in the context of applying the IBP relations was proposed in Baikov:2005nv ; Baikov:2007zza and was successfully used in the number of calculations (see, for example Baikov:2008jh ; Baikov:2009bg ; Baikov:2010je ; Baikov:2012er ; Baikov:2012zm ; Baikov:2012zn ; Baikov:2014qja ; Baikov:2016tgj ; Baikov:2017ujl ). From a technical point of view this requires expansion in the limit $1/d\rightarrow 0$ of integrals with integrands depending on $P(x_{i})^{d/2}$ where $P(x_{i})$ is some multivariate polinomial. This expansion causes the size of intermediate expressions to rapidly increase and calculations to a sufficiently high order of $1/d$ become possible only after manual optimization (smart redefinition of variables, etc.), which was carried out individually for different subclasses of integrals. Bearing in mind the significant increase in the subclasses number at the next loop(s) level, in this article we reformulate the problem as constructing solutions to IBP relations in the form of formal series of $1/d$ and propose linear substitutions that transform these relations into an explicitly recursive form.
The plan of the paper is the following. In section 2 we fix the notations and briefly recall the main points of the Baikov:1996rk ; Baikov:1996iu ; Baikov:2005nv ; Baikov:2007zza . Then in section 3 we introduce (for one-parameter case) specific linear substitutions, which we apply in section 4 to the IBP relations. In the next section 5 we consider several examples of massless (with one massive line) vacuum integral families up to the 7-loop level. Finally, in section 6 we discuss a possible extension to more complex cases.
2 Integration by parts relations for Feynman integrals
Let us consider the $d$-dimensional $L$-loop vacuum Feynman integrals:
$F(n_{1},\ldots,n_{N};d;m^{2})=\int\ldots\int\frac{\mbox{d}^{d}p_{1}\ldots\mbox{d}^{d}p_{L}}{D_{1}^{n_{1}}\ldots D_{N}^{n_{N}}}\equiv(m^{2})^{Ld/2-\Sigma n_{i}}f(n_{1},\ldots,n_{N};d),$ (1)
where $p_{i}$ ($i=1,\ldots,L$}) are loop momenta and
$D_{a}$ $=$ $A^{ij}_{a}p_{i}\cdot p_{j}-\mu_{a}m^{2},\quad a=1,\ldots,N=L(L+1)/2$ (2)
are inverse propagators. The matrix $A^{ij}_{a}$ is supposed to be symmetrical with respect to upper indices. The $1/D_{a}$ can be raised as to positive powers (“denominators” or “lines”) as non-positive (“numerators”). Summation over repeating indices is implied and in many cases (unless it’s confusing) it will be omitted. In what follows we will set $m^{2}=1$; we will also omit $d$ in the arguments and use the shorthand $(n_{1},\ldots,n_{N})\equiv(\underline{n})$.
We assume that the set (2) is complete and we can express scalar products $p_{i}\cdot p_{j}$ through $D_{a}$:
$p_{i}\cdot p_{k}$ $=$ $\hat{A}^{ik}_{a}(D_{a}+\mu_{a})\,.$ (3)
The matrix $\hat{A}^{ij}_{a}$ is also symmetrical with respect to $(ij)$ and is inverse to $A^{kl}_{a}$ in the sense that
$\hat{A}^{ij}_{a}A^{kl}_{a}$ $=$ $(\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk})/2,\quad A^{ik}_{a}\hat{A}^{ik}_{b}=\delta_{ab}.$ (4)
Please note that in Baikov:2005nv for the coefficients in (3) we used the notation $\tilde{A}$; we change it to $\hat{A}$ because we want to reserve the $\sim$” sign for “soft” objects (see section 4).
Let’s define
${\bf I}^{-}_{a}f(\ldots,n_{a},\ldots)$ $\equiv$ $f(\ldots,n_{a}-1,\ldots),$
${\bf I}^{+}_{a}f(\ldots,n_{a},\ldots)$ $\equiv$ $n_{a}f(\ldots,n_{a}+1,\ldots)$
(no summation in $n_{a}$ in the last relation).
IBP relations can be obtained Chetyrkin:1981qh ; Tkachov:1981wb by inserting the operator $(\partial/\partial p_{i})\cdot p_{k}$ in the integrand, equating the integral of the derivative to zero and expressing all the terms resulting from the differentiation through the integrals with shifted $(\underline{n})$. Starting from eq. (1) we obtain (up to $1/2$ factor)
$0=R_{ik}(\underline{n})$ $=$ $\big((d/2)\,\delta_{ik}-\hat{A}^{kl}_{a}({\bf I}^{-}_{a}+\mu_{a})A_{b}^{il}{\bf I}_{b}^{+}\big)\,f(\underline{n})\,$, (5)
$=$ $\big((d-L-1)/2\,\delta_{ik}-A_{b}^{il}{\bf I}_{b}^{+}\,\hat{A}^{kl}_{a}({\bf I}^{-}_{a}+\mu_{a})\big)f(\underline{n})\,$, (6)
where we use ${\bf I}^{-}_{a},{\bf I}^{+}_{b}]=\delta_{ab}$ to derive eq. (6).
[…]
In many practically important cases eq. (5) can be used to relate integrals with arbitrary values of $(\underline{n})$ to a linear combination of a finite number of irreducible (master) integrals:
$f(\underline{n})$ $=$ $\sum_{k}c_{k}(\underline{n})\,M_{k},\quad M_{k}=f(\underline{n}_{k}),$ (7)
where $f(\underline{n}_{k})$ are chosen as the most convenient for direct calculation, usually with values $(\underline{n}_{k})$ from the set (1,0,+1)(-1,0,+1).
Such a reduction can be achieved as by deriving relations with an explicitly recursive structure (MINCER Gorishnii:1989gt for 3-loop propagator integrals, Forcer Ruijl:2017cxj for the next loop level) as by Laporta approach Laporta:2000dsw with a number of implementations Anastasiou:2004vj ; Smirnov:2019qkx ; Klappert:2020nbg ; Studerus:2009ye ; vonManteuffel:2012np . In any case, reduction to irreducible integrals during actual calculations requires huge computer resources, so any progress in this field is very desirable. In this context, a direct calculation of c(n) was proposed Baikov:1996rk ; Baikov:1996iu . Indeed, if $M_{k}$ are linearly independent, then $c_{k}(\underline{n})$ are rational in $d$ solutions of IBP relations and can be obtained from any other solutions set as linear combinations satisfying the condition $c_{i}(\underline{n}_{k})=\delta_{ik}$.
There are natural boundary conditions on $c_{k}(\underline{n})$, which follow from the assumption that no additional denominators (lines) appear on the right side of eq. (7) (since the purpose of the reduction procedure is to relate the original integral with simpler ones):
$c_{k}(\underline{n})$ $\neq$ $0\quad\mbox{only if}\,\,n_{a}>0\,\,\mbox{for all}\,\,a\in H_{k},$
where $H_{k}=\{a_{1},\dots,a_{N_{k}}\}$ is the set of line numbers of the irreducible integral $M_{k}$ and $N_{k}$ is the total number of lines in $M_{k}$. In some cases we will also use an alternative boundary condition for $a\notin H_{k}$:
$f(\underline{n})$ $\neq$ $0\quad\mbox{only if}\,\,n_{a}<1\,\,\mbox{for all}\,\,a\notin H_{k}.$ (8)
Here and below we consider $f(\underline{n})$ as some solution of recurrence relations (and integrals (1) as a particular example).
In Baikov:1996rk ; Baikov:1996iu $c_{k}(\underline{n})$ were constructed using the auxiliary integral representation
$f(\underline{n})$ $=$ $\int dx_{1}\dots dx_{N}\,f(\underline{x})/(x_{1}^{n_{1}}..\ x_{N}^{n_{N}}),$ (9)
where $f(\underline{x})$ should be chosen according to recurrence relations; for example, relations for integrals (1) require
$f(\underline{x})=\det(\hat{A}^{ik}_{a}(x_{a}+\mu_{a}))^{(d-L-1)/2}.$ (10)
Integrals (9) can be expanded in the $1/d\rightarrow 0$ limit Baikov:2005nv ; Baikov:2007zza but as mentioned in the introduction this is difficult to extend to the next loop(s) level. Therefore, as a more promising alternative, we are going to explore ways to find the solutions to IBP relations as formal series in $1/d$.
3 Linear substitutions in recurrence relations: one-parameter case
Let’s consider the one-parameter recurrence relation:
$0$ $=$ $R(n)=R({\bf I}_{f}^{-},{\bf I}_{f}^{+})f(n),$ (11)
where by definition
${\bf I}_{f}^{+}f(n)$ $\equiv$ $nf(n+1),\quad{\bf I}_{f}^{-}f(n)\equiv f(n-1),$ (12)
and hence
$\big[{\bf I}_{f}^{-},{\bf I}_{f}^{+}\big]$ $\equiv$ ${\bf I}_{f}^{-}{\bf I}_{f}^{+}-{\bf I}_{f}^{+}{\bf I}_{f}^{-}=1,$
$\big[{\bf I}_{f}^{-},({\bf I}_{f}^{+})^{k}\big]$ $=$ $k\,({\bf I}_{f}^{+})^{k-1},\quad\big[{\bf I}_{f}^{+},({\bf I}_{f}^{-})^{k}\big]=-k\,({\bf I}_{f}^{-})^{k-1},$
$\big[{\bf I}_{f}^{-},S({\bf I}_{f}^{+})\big]$ $=$ $\partial_{{\bf I}_{f}^{+}}S({\bf I}_{f}^{+}),\quad\big[{\bf I}_{f}^{+},S({\bf I}_{f}^{-})\big]=-\partial_{{\bf I}_{f}^{-}}S({\bf I}_{f}^{-}),$
where $S({\bf I})$ is any formal series in $\bf I$}.
We are going to study the transformation of eq. (11) after linear redefinition of $f(n)$. The idea is to replace $f(n)$ with some combination of $g(n)$ to obtain a simpler recurrence relation for $g(n)$, solve it and re-express the results in terms of $f(n)$.
Let’s start with eq. (11) with the boundary condition
$f(n<1)=0.$ (13)
Due to eq. (13) if k>n1k>n-1 then $({\bf I}_{f}^{-})^{k}f(n)=f(n-k)=0$, so
$g(n)$ $=$ $e^{\lambda{\bf I}_{f}^{-}}f(n)\equiv\sum_{k=0}^{\infty}\frac{(\lambda{\bf I}_{f}^{-})^{k}}{k!}f(n)$ (14)
will be equal to the finite linear combination of $f(k)$ with $k=1,\dots,n$. We are going to derive the recurrence relation for $g(n)$, which follows from eq. (11). To do this we need to find the relations between operators ${\bf I}_{f}^{\pm}$ defined by eq. (12) and ${\bf I}_{g}^{\pm}$, acting in the same way on $g(n)$. Let’s start with ${\bf I}_{g}^{+}$:
${\bf I}_{g}^{+}g(n)$ $\equiv$ $ng(n+1)=ne^{\lambda{\bf I}_{f}^{-}}f(n+1)=e^{\lambda{\bf I}_{f}^{-}}nf(n+1)=e^{\lambda{\bf I}_{f}^{-}}{\bf I}_{f}^{+}f(n)$
$=$ $\big(\big[e^{\lambda{\bf I}_{f}^{-}},{\bf I}_{f}^{+}\big]+{\bf I}_{f}^{+}e^{\lambda{\bf I}_{f}^{-}}\big)f(n)=\big(\partial_{{\bf I}_{f}^{-}}(e^{\lambda{\bf I}_{f}^{-}})+{\bf I}_{f}^{+}e^{\lambda{\bf I}_{f}^{-}}\big)f(n)=(\lambda+{\bf I}_{f}^{+})e^{\lambda{\bf I}_{f}^{-}}f(n)$
$=$ $({\bf I}_{f}^{+}+\lambda)g(n).$
The relation for ${\bf I}_{g}^{-}$ is more simple:
${\bf I}_{g}^{-}g(n)$ $\equiv$ $g(n-1)=e^{\lambda{\bf I}_{f}^{-}}f(n-1)=e^{\lambda{\bf I}_{f}^{-}}{\bf I}_{f}^{-}f(n)={\bf I}_{f}^{-}e^{\lambda{\bf I}_{f}^{-}}f(n)$
$=$ ${\bf I}_{f}^{-}g(n).$
Now we can invert eq. (14):
$f(n)$ $=$ $e^{-\lambda{\bf I}_{f}^{-}}g(n)$ (15)
$=$ $e^{-\lambda{\bf I}_{g}^{-}}g(n).$
Again, due to eqs. (1314) it turns out that $g(n<1)=0$ and eq. (15) consists of a finite number of terms.
The recurrence relation for $g(n)$ follows from the linear combination of eqs. (11) with shifted $n$ (${\bf I}_{R}^{-}R(n)\equiv R(n-1)$):
$0$ $=$ $e^{\lambda{\bf I}_{R}^{-}}R(n)=R({\bf I}_{f}^{-},{\bf I}_{f}^{+})e^{\lambda{\bf I}_{f}^{-}}f(n)=R({\bf I}_{f}^{-},{\bf I}_{f}^{+})g(n)$
$=$ $R({\bf I}_{g}^{-},{\bf I}_{g}^{+}-\lambda)g(n).$
The case with the boundary condition
$f(n>0)=0$ (16)
is considered similarly, we present the final formulas:
$g(n)$ $=$ $e^{\lambda{\bf I}_{f}^{+}}f(n),\quad g(n>0)=0,$
$f(n)$ $=$ $e^{-\lambda{\bf I}_{g}^{+}}g(n),$ (17)
${\bf I}_{g}^{-}$ $=$ ${\bf I}_{f}^{-}-\lambda,\quad{\bf I}_{g}^{+}={\bf I}_{f}^{+},$
$0$ $=$ $R({\bf I}_{g}^{-}+\lambda,{\bf I}_{g}^{+})g(n).$ (18)
The boundary condition (16) corresponds to the condition (8), that is, the case when the master integral $M_{k}$ does not have the line $D_{a}$. However, there may be such a line on the left side of eq. (7), which means that $n_{a}$ in $c_{k}(\underline{n})$ can be positive. So, let us extend the substitution (17) to positive $n$. For $n<1$ we can rewrite $g(n)=({\bf I}_{g}^{-})^{-n}g(0)$ and then eq. (17) as follows:
$f(n)$ $=$ $e^{-\lambda{\bf I}_{g}^{+}}({\bf I}_{g}^{-})^{-n}g(0)=({\bf I}_{g}^{-}+\lambda)e^{-\lambda{\bf I}_{g}^{+}}({\bf I}_{g}^{-})^{-n-1}g(0)=\dots=({\bf I}_{g}^{-}+\lambda)^{-n}e^{-\lambda{\bf I}_{g}^{+}}g(0)$ (19)
$=$ $({\bf I}_{g}^{-}+\lambda)^{-n}g(0).$
For $f(n<1)$ the binomial expansion reproduces the finite sum (17) as it should. For $f(n>0)$ we expand the right side of eq. (19) into a formal series in ${\bf I}_{g}^{-}$. From the negative binomial series we get
$f(n>0)$ $=$ $({\bf I}_{g}^{-}+\lambda)^{-n}g(0)$ (20)
$=$ $\sum_{k=0}^{\infty}(-1)^{k}\frac{(n+k-1)!}{(n-1)!\,k!}\lambda^{-n-k}({\bf I}_{g}^{-})^{k}g(0)$
$=$ $\sum_{k=0}^{\infty}(-1)^{k}\frac{(n+k-1)!}{(n-1)!\,k!}\lambda^{-n-k}g(-k),$
that is the infinite sum of $g(n<1)$. However, as we will see in section 4 one can choose $\lambda$ so that the $k$-term in the sum (20) has leading asymptotics $d^{-k/2}$ (see eq. (21)), therefore, by keeping enough terms in this sum and assuming that $g(n<1)$ satisfies eq. (18) we can satisfy (11) in any desired order in $1/d$. As a result, $f(n)$ can be parameterized using $g(k<1)$ for any value of nn. [^1] Note that a direct extension of eq. (17) for $n>0$ produces an infinite sum over $k$ of terms with undesirable leading asymptotic $d^{k/2}$ (see eq. (21)). In what follows, by the expression (17) in the case (16) we will mean the extension (20) for $n>0$ (in the $\lambda=0$ case when (20) is not defined we will set $f(n)=g(n)$).
Finally, if we rescale $f(n)=\alpha^{n}g(n)$, then eq. (11) transforms into
$0$ $=$ $R({\alpha^{-1}\bf I}_{g}^{-},\alpha{\bf I}_{g}^{+})g(n).$
4 Linear substitutions in recurrence relations: Feynman integrals
Let us now return to the general case and construct, if possible, a solution to eq. (5) for some set of propagator indices $H_{k}=\{a_{1},\dots,a_{N_{k}}\}$. We will call the index $a\in H_{k}$ “hard”. (substitutions (15) will be applied), and $a\notin H_{k}$ “soft’ (substitutions (1720)).
For any indexed object $x_{a}$ we define projections onto the “hard” and “soft” subspaces: ${\overline{x}}_{a}\equiv x_{\overline{a}}=$ (if $a\in H_{k}$ then $x_{a}$ else 0) and ${\tilde{x}}_{a}\equiv x_{\tilde{a}}=x_{a}-{\overline{x}}_{a}$; as the result $x_{a}=\overline{x}_{a}+\tilde{x}_{a}$. For $x_{a}={\bf I}_{a}^{\pm}$ we have ${\bf I}_{a}^{\pm}={\tilde{\bf I}}_{a}^{\pm}+{\overline{\bf I}}_{a}^{\pm},\,\big[{\tilde{\bf I}}_{a}^{\pm},{\overline{\bf I}}_{b}^{\pm}\big]=0.$
Let us redefine
$d/2$ $=$ $1/\alpha^{2},\quad\alpha\rightarrow 0.$
According to section 3 the substitutions
$f(\underline{n})$ $=$ $\exp(\alpha^{-2}\overline{\lambda}_{a}\,\overline{{\bf I}}_{f^{\prime},a}^{-}-\tilde{\lambda}_{a}\tilde{\bf I}_{f^{\prime},a}^{+})f^{\prime}(\underline{n}),$
$f^{\prime}(\underline{n})$ $=$ $\alpha^{N_{k}-\sum_{a}n_{a}}g(\underline{n})$ (21)
will change
$\overline{{\bf I}}_{f,a}^{+}=\alpha^{-2}(\alpha\overline{{\bf I}}_{g,a}^{+}+\overline{\lambda}_{a}),\quad\tilde{{\bf I}}_{f,a}^{-}=\alpha\tilde{{\bf I}}_{g,a}^{-}+\tilde{\lambda}_{a},\quad\tilde{I}_{f}^{+}=\alpha^{-1}\tilde{I}_{g}^{+},\quad\overline{I}_{f}^{-}=\alpha\overline{I}_{g}^{-},$
and therefore eq. (5) (multiplied by $\alpha^{2}$) will be transformed into
$0$ $=$ $\Big(\delta_{ik}-\hat{A}^{kl}(\alpha\,{\bf I}^{-}_{g}+\tilde{\lambda}+\mu)A^{li}\big({\alpha\,\bf I}_{g}^{+}+\overline{\lambda}\big)\Big)\,g(\underline{n})\,$ (22)
$=$ $\Big(\delta_{ik}-\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}\,\overline{\lambda}$
$-\alpha\,\big(\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}{\bf I}_{g}^{+}+\hat{A}^{kl}{\bf I}^{-}_{g}A^{li}\overline{\lambda}\big)-\alpha^{2}\hat{A}^{kl}{\bf I}^{-}_{g}A^{li}{\bf I}_{g}^{+}\Big)g(\underline{n}).$
Here and in some places below we omit the “propagator” indices over which the summation is performed, assuming that matrices $A,\hat{A}$ without a subscript are summed with the object located on the right. Also, compared to eq. (5), we have changed the order of the upper indices of the matrix $A^{kl}_{a}$ (since it is symmetrical).
Eqs. (22) generate relations for the expansion coefficients $g_{j}(\underline{n})$ in $g(\underline{n})=\sum_{j=0}^{\infty}\alpha^{j}\ g_{j}(\underline{n})$:
$0$ $=$ $\big(\delta_{ik}-\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}\overline{\lambda}\big)g_{j}(\underline{n})$ (23)
$-\big(\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}{\bf I}^{+}+\hat{A}^{kl}{\bf I}^{-}A^{li}\overline{\lambda}\big)g_{j-1}(\underline{n})$
$-\hat{A}^{kl}{\bf I}^{-}A^{li}{\bf I}^{+}g_{j-2}(\underline{n}).$
From here on we omit the indices $f,f^{\prime},g$ in $\bf I$$^{\pm}$ if the “default” action is implied.
Equations for $g_{0}(\underline{n})$ do not contain ${\bf I}_{a}^{\pm}$ operators and in general lead to $g_{0}(\underline{n})=0$ (and as a result $g_{j}(\underline{n})=0$), therefore, to obtain non-trivial solutions it is necessary to choose $\lambda_{a}$ satisfies the equations
$\delta_{ik}$ $=$ $\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}\,\overline{\lambda}.$ (24)
Eqs. (23) with the condition (24) become simpler:
$0$ $=$ $\big(\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}{\bf I}^{+}+\hat{A}^{kl}{\bf I}^{-}A^{li}\overline{\lambda}\big)g_{j}(\underline{n})$ (25)
$+\hat{A}^{kl}{\bf I}^{-}A^{li}{\bf I}^{+}g_{j-1}(\underline{n}).$
We are going to use eqs. (25) to reduce $g_{j}(\underline{n})$ (up to $g_{j-1}(\underline{n})$ terms) to “corner” values: “hard” $\overline{n}_{a}\to 1$ and “soft” $\tilde{n}_{a}\to 0$ (assuming “soft” $\tilde{\lambda}_{a}\neq 0$, see the end of section 3). This would be possible if the $g_{j}(\underline{n})$ part of eqs. (25) could be solved explicitly with respect to the operators ($\overline{\bf I}_{a}^{+},\tilde{\bf I}_{a}^{-}$), increasing the distance from the corner. Let us multiply eqs. (25) by $(\hat{A}_{a}A\overline{\lambda})^{ik}$:
$0$ $=$ $\Big(\mbox{Tr}\big(\hat{A}_{a}A\overline{\lambda}\,\hat{A}(\tilde{\lambda}+\mu)A{\bf I}^{+}\big)+\mbox{Tr}\big(\hat{A}_{a}A\overline{\lambda}\,\hat{A}{\bf I}^{-}A\overline{\lambda}\big)\Big)g_{j}(\underline{n})+\mbox{Tr}(\dots)g_{j-1}(\underline{n})$ (26)
$=$ $\big({\bf I}_{a}^{+}+\mbox{Tr}(\hat{A}_{a}A\overline{\lambda}\,\hat{A}_{b}A\overline{\lambda}){\bf I}_{b}^{-}\big)g_{j}(\underline{n})+\mbox{Tr}(\hat{A}_{a}A\overline{\lambda}\,\hat{A}{\bf I}^{-}A{\bf I}^{+})g_{j-1}(\underline{n}).$
Here we omit the repeated “loop” indices, $\mbox{Tr}(\cdot)$ means the contraction of outer indices $(ik)$; we also use the condition (24) and $\mbox{Tr}(\hat{A}_{a}A_{b})=\delta_{ab}$. In the case of “soft” $\tilde{a}$, the $g_{j}(\underline{n})$ part of eqs. (26) depends on increasing $\tilde{\bf I}_{a}^{-}$ and decreasing $(\tilde{\bf I}_{a}^{+},\,\overline{\bf I}_{a}^{-})$ operators and if
$\det_{\tilde{a}\tilde{b}}\mbox{Tr}(\hat{A}_{\tilde{a}}A\overline{\lambda}\,\hat{A}_{\tilde{b}}A\overline{\lambda})$ $\neq$ $0,$ (27)
where both indices $a,b$ are “soft”, we can express increasing “soft” operators $\tilde{\bf I}_{a}^{-}$ in terms of decreasing ones:
$\tilde{\bf I}_{a}^{-}$ $\rightarrow$ $(\tilde{\bf I}_{a}^{+},\,\overline{\bf I}_{a}^{-}).$ (28)
For “hard” $\overline{a}$ eqs. (26) explicitly express the remaining increasing operators $\overline{\bf I}_{a}^{+}$ in terms of $(\tilde{\bf I}_{a}^{-},\,\overline{\bf I}_{a}^{-})$ and due to eqs. (28) in terms of decreasing $(\tilde{\bf I}_{a}^{+},\,\overline{\bf I}_{a}^{-})$ only, so as the result we have an explicit reduction to “corner” values
$(\tilde{\bf I}_{a}^{-},\,\overline{\bf I}_{a}^{+})$ $\rightarrow$ $(\tilde{\bf I}_{a}^{+},\,\overline{\bf I}_{a}^{-}).$ (29)
The $L(L+1)/2$ equations (29) are linear combinations (by means of eqs. (26)) of the $L^{2}$ eqs. (25), so we can expect additional linear relations on ${\bf I}_{a}^{\pm}$, which can lead to a contradiction. Fortunately if (27) and hence (29) are valid this is not the case. Indeed, eqs. (26) are the combinations of eqs. (29). Then the $g_{j}(\underline{n})$ parts of eqs. (25) are the linear combinations of the $g_{j}(\underline{n})$ parts of eqs. (26) with the coefficients $(\hat{A}(\tilde{\lambda}+\mu)A_{a})^{ki}$, so substituting (29) into eqs. (25) nullifies this part and no additional relations appear.
The case when some of the “soft” $\tilde{\lambda}_{a}=0$ is more complicated, since the corresponding $\tilde{n}_{a}$ can be both negative and positive and finding out the possibility of reducing them to a minimum value requires a more detailed study of the $g_{j}(\underline{n})$ parts of eqs. (25).
Eqs. (24) lead to several useful consequences. First, $\mbox{Tr}(\cdot)$ of eqs. (24) due to eq. (4) leads to
$L$ $=$ $\overline{\mu}_{a}\,\overline{\lambda}_{a}.$ (30)
Then, $\det(\cdot)$ of eqs. (24) leads to
$1$ $=$ $\det\big(\hat{A}(\tilde{\lambda}+\mu)\big)\det(A\overline{\lambda})$ (31)
$\Rightarrow$ $\det\big(\hat{A}(\tilde{\lambda}+\mu)\big)\neq 0.$
Next, $\hat{A}(\tilde{\lambda}+\mu)$ and $A\overline{\lambda}$ considered as matrices with indices $(ik)$ are reciprocal
$A\overline{\lambda}$ $=$ $\hat{A}(\tilde{\lambda}+\mu)^{-1}.$ (32)
Finally, let us multiply eqs. (24) by $\big(\hat{A}_{a}\hat{A}(\tilde{\lambda}+\mu)^{\mbox{\small Adj}}\big)^{ik}$:
$\partial_{a}\det\big(\hat{A}(\tilde{\lambda}+\mu)\big)$ $=$ $\overline{\lambda}_{a}\,\det\big(\hat{A}(\tilde{\lambda}+\mu)\big),$ (33)
where $A^{\mbox{\small Adj}}\equiv\det(A)\,A^{-1}$, $\partial_{a}\det(A\lambda)\equiv\partial_{x_{a}}\det(A^{ik}_{a}x_{a})|_{x_{a}=\lambda_{a}}$. For “soft” $\tilde{a}$ eqs. (33) give
$\partial_{\tilde{a}}\det(\hat{A}(\tilde{\lambda}+\mu))$ $=$ $0.$ (34)
Using eqs. (32) we can re-express eq. (27) via $\tilde{\lambda}_{a}$:
$0$ $\neq$ $\det_{\tilde{a}\tilde{b}}\mbox{Tr}\big(\hat{A}_{\tilde{a}}A\overline{\lambda}\,\hat{A}_{\tilde{b}}A\overline{\lambda}\big)$ (35)
$=$ $\det_{\tilde{a}\tilde{b}}\Big(\mbox{Tr}\big(\hat{A}_{\tilde{a}}\hat{A}(\tilde{\lambda}+\mu)^{-1}\hat{A}_{\tilde{b}}\hat{A}(\tilde{\lambda}+\mu)^{-1}\big)\Big)$
$\propto$ $\det_{\tilde{a}\tilde{b}}\Big(\partial_{\tilde{a}}\partial_{\tilde{b}}\det\big(\hat{A}(\tilde{\lambda}+\mu)\big)/\det\big(\hat{A}(\tilde{\lambda}+\mu)\big)\Big),$
where in the third line we use eqs. (34) and standard matrix formulas
$\partial_{x}\det(A)$ $=$ $\det(A)\,\mbox{Tr}\big(A^{-1}\partial_{x}A\big),\quad\partial_{x}(A^{-1})=-A^{-1}\,\partial_{x}(A)\,A^{-1}.$
Please note that eqs. (34, 31, 35) correspond to eqs. (16) from Baikov:2005nv , that is to the conditions for the existence of the $1/d$ expansion of the integral representation (9, 10).
5 Examples
Let us consider 2-loop example with inverse propagators
$D_{1}=p_{1}^{2},\quad D_{2}=p_{2}^{2},\quad D_{3}=(p_{1}+p_{2})^{2}-1.$
Recurrence relations (5) are
$R_{11}(\underline{n})$ $=$ $1/2(d-2\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}-\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{+})f(\underline{n}),$
$R_{12}(\underline{n})$ $=$ $1/2(\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}+\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{1}^{+}-\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{1}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{1}^{+}-\overline{\bf I}_{3}^{+})f(\underline{n}),$
$R_{21}(\underline{n})$ $=$ $1/2(\overline{\bf I}_{1}^{-}\overline{\bf I}_{2}^{+}-\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{2}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{2}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{2}^{+}-\overline{\bf I}_{3}^{+})f(\underline{n}),$
$R_{22}(\underline{n})$ $=$ $1/2(d+\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}-2\overline{\bf I}_{2}^{-}\overline{\bf I}_{2}^{+}-\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{+})f(\underline{n}).$ (36)
Let us assume that all indices are “hard”,
[…]
and substitute (remember that $d/2=\alpha^{-2}$):
$f(\underline{n})$ $=$ $\exp(\alpha^{-2}\sum_{a=1}^{3}\overline{\lambda}_{a}\overline{\bf I}_{a}^{-})f^{\prime}(\underline{n}),\quad f^{\prime}(\underline{n})=\alpha^{3-\sum_{a=1}^{3}n_{a}}g(\underline{n}),$ (38)
$f^{\prime}(\underline{n}),\,g(\underline{n})$ $=$ $0\,\,\,\mbox{if any of}\,\,(n_{1},n_{2},n_{3})<1.$ (39)
These substitutions correspond to
$\overline{{\bf I}}_{f,a}^{+}=\alpha^{-1}\overline{{\bf I}}_{g,a}^{+}+\alpha^{-2}\overline{\lambda}_{a},\quad\overline{{\bf I}}_{f,a}^{-}=\alpha\overline{{\bf I}}_{g,a}^{-}.$
The $\alpha^{-2}$ parts of (transformed) eqs. (36) lead to equations on $\overline{\lambda}_{a}$ with the solution
$\overline{\lambda}_{1}=-2,\quad\overline{\lambda}_{2}=-2,\quad\overline{\lambda}_{3}=2,$ (40)
while the $\alpha^{-1}$ and $\alpha^{0}$ parts of the equations $\{0=R_{11},0=R_{12},0=R_{21}\}$ (after substituting eqs. (40) and multiplying by $\alpha$) lead to recursive relations
$\overline{\bf I}_{1}^{+}g(\underline{n})$ $=$ $\big(-2\overline{\bf I}_{1}^{-}-6\overline{\bf I}_{2}^{-}+2\overline{\bf I}_{3}^{-}+\alpha\,(3\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}+2\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{1}^{+}-2\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{1}^{+})\big)g(\underline{n}),$
$\overline{\bf I}_{2}^{+}g(\underline{n})$ $=$ $\big(-6\overline{\bf I}_{1}^{-}-2\overline{\bf I}_{2}^{-}+2\overline{\bf I}_{3}^{-}+\alpha\,(2\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}+\overline{\bf I}_{1}^{-}\overline{\bf I}_{2}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{2}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{2}^{+})\big)g(\underline{n}),$
$\overline{\bf I}_{3}^{+}g(\underline{n})$ $=$ $\big(2\overline{\bf I}_{1}^{-}+2\overline{\bf I}_{2}^{-}-2\overline{\bf I}_{3}^{-}+\alpha\,(-2\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}-\overline{\bf I}_{1}^{-}\overline{\bf I}_{3}^{+}+\overline{\bf I}_{2}^{-}\overline{\bf I}_{3}^{+}-\overline{\bf I}_{3}^{-}\overline{\bf I}_{3}^{+})\big)g(\underline{n}).$ (41)
Eqs. (38-41) allow us to derive a relation between $f(\underline{n})$ for a particular $(\underline{n})$ and $f(1,1,1)$; as an example, let us do this for $f(2,1,1)$. In the first step we need to express $f(2,1,1)$ and $f(1,1,1)$ in terms of $g(\underline{n})$. According to eq. (38) with (40)
$f(2,1,1)$ $=$ $(1-2\alpha^{-2}\overline{\bf I}_{1}^{-}+\dots)f^{\prime}(2,1,1)$ (42)
$=$ $f^{\prime}(2,1,1)-2\alpha^{-2}f^{\prime}(1,1,1)$
$=$ $\alpha^{-1}g(2,1,1)-2\alpha^{-2}g(1,1,1),$
$f(1,1,1)$ $=$ $g(1,1,1).$ (43)
From here on, dots denote terms that do not contribute due to the conditions (37, 39). In the second step we reduce the right-hand side of eq. (42) to $g(1,1,1)$. The first equation (41) for $(\underline{n})=(1,1,1)$ gives
$g(2,1,1)=(3\alpha\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}+\dots)g(1,1,1)=3\alpha g(1,1,1),$ (44)
and hence
$f(2,1,1)=(3-2\alpha^{-2})g(1,1,1)=(3-d)g(1,1,1).$ (45)
From eqs. (43, 45) we finally obtain
f(2,1,1)=(3d)f(1,1,1).f(2,1,1)=(3-d)f(1,1,1). (46)
Please note that in this example the recursion stopped due to condition (39), so the result (in terms of $\alpha$) is exact; in general one should stop when the desired order of $\alpha$ is reached (for compactness we do not expand $g(\underline{n})$ into a series in $\alpha$ explicitly). Also note that only even powers of $\alpha$ (integer powers of $d$) contribute to the final relation (46), unlike the relations (42, 44).
The standard Laporta’s approach, that is solving eqs. (36, 37) for some (sufficiently large) set of values for $(\underline{n})$, gives
$0$ $=$ $R_{11}(1,1,1)-R_{21}(1,1,1)$
$=$ $1/2(d-2\overline{\bf I}_{1}^{-}\overline{\bf I}_{1}^{+}-\overline{\bf I}_{2}^{-}\overline{\bf I}_{2}^{+}+\overline{\bf I}_{2}^{+}+\dots)f(1,1,1)$
$=$ $(d/2-3/2)f(1,1,1)+1/2f(2,1,1),$
which confirms eq. (46).
Calculating the parameters $\lambda_{a}$ is the key point for obtaining recursive relations for coefficients of $1/d$ series. To find $\lambda_{a}$, it is necessary to solve eqs. (24), that is, a system of quadratic equations for $L(L+1)/2$ variables (for example, 28 variables for the 7-loop case). However, the bilinear structure of eqs. (24) allows us to significantly reduce this number. First, using the linear equation (30) we can exclude one of $\lambda_{a}$. Moreover, we can obtain additional linear relations on $\lambda_{a}$ checking general combinations of eqs. (24) with polynomials in $\lambda_{a}$ as coefficients. If such relations are found, with their help we can eliminate some $\lambda_{a}$ from eqs. (24) and repeat the same steps for the resulting equations.
Figure 1 presents “hard” $\overline{\lambda}_{a}$ (“soft” $\tilde{\lambda}_{a}$ can be recovered using eqs. (32)) for leading sectors of some integral families up to the 7-loop order. Polynomials of the first degree as coefficients were sufficient to reduce the original system of equations to an equation for one variable. [^2] Symbolic manipulations were performed using REDUCE FreeREDUCE ).
L=5L=5cceeeecceeeebbbbaabbbbcc
L=6L=6cceeeeeecceeeeeebbbbaabbbbcccc
L=7L=7cceeeeeeeecceeeeeeeebbbbaabbbbcccccc
L=6L=6eeeeeeeeeeeeccccccccbbbbaabbbb
L=7L=7eeeeeeeeeeeeeeeeccccccccccbbbbaabbbb
Figure 1: $\overline{\lambda}_{a}$ (roots of eqs. (24)) for some vacuum diagrams. Ordinary lines are massless, double line is massive with $\mu=1$; $a_{L}=L$, $b_{L}=-\frac{L-2}{2}e_{L}$, $c_{L}=\frac{L(L-2)}{2}e_{L}-L(L-1)$;
for planar diagrams (first row) $e_{5}=\frac{20}{9}$, $e_{6}=\frac{15}{8}$, $e_{7}=(\pm 28\,(61)^{1/2}+1232)/625$;
for nonplanar (second row) $e_{6}=(\pm 3\,(41)^{1/2}+57)/16$, $e_{7}=(\pm 168\,(26)^{1/2}+2268)/625$.
The results show that in the 5-, 6-loop planar examples there is at least one variant of the substitutions which transform IBP relations to a recursive form, in the 7-loop planar and in the 6-, 7-loop nonplanar cases there are at least two. The $L$-dependence of the results was obtained empirically, but there are some indications of its validity for an arbitrary order, we hope to clarify this point in future publications.
Please note that eqs. (16) from Baikov:2005nv (eqs. (34) in the notation of this article) at the 7-loop level lead to polynomial equations of the 6th degree for 10 variables, which is a much more difficult task.
6 Conclusion
In this paper we present a method to construct solutions of IBP relations for Feynman integrals as formal series in the limit $1/d\rightarrow 0$. The idea is to use substitutions of the form (1517) which transform IBP relations into an explicitly recursive form. From a technical point of view, it is necessary to calculate the constants $\lambda$ as solutions to a system of bilinear quadratic equations (24). Although this is relatively simple (and therefore can be automated) in the case where one line is massive and the others are massless, more general kinematic configurations may be more difficult to study. In particular, it may happen that the number of $1/d$ solutions constructed will be less than the number of master integrals and, therefore, not enough to calculate all $c_{i}(\underline{n})$ coefficients in eq. (7). However, even in this case they provide some linear relations between $c_{i}(\underline{n})$ that can be used (in combination with other data) for their reconstruction or at least for verification.
Please note that although in this article we are dealing with recurrence relations for vacuum integrals, the substitutions (15, 17) can also be applied to the case with external momenta. In this context, it is worth mentioning the work Mizera:2019vvs , where the large $d$ limit was used to obtain the differential equations for master integrals. We hope that studying possible transformations of the corresponding IBP relations will help in the difficult task of evaluating “multi-legs” integrals; we leave this for future research.